what is the length of each side of a rhombus with the perimeter of 44 inches?
given the information can you determine the lengths?

Answer by jim_thompson5910(35256)

A rhombus is a quadrilateral with 4 equal sides. So this means that if x is the length of one side, then all sides are equal to x.


The perimeter is then x+x+x+x = 4x

Because we know that the perimeter is really 44, we can say 4x = 44


Solve for x to get: x = 44/4 = 11


So the length of each side is 11 inches.